# Dysca: A Dynamic and Scalable Benchmark for Evaluating Perception Ability of LVLMs

### v2

@@ B.1 General Discussion
Synthesis Data for Training / Fine-tuning. The use of synthetic data for model training / fine-tuning has been adopted in the field of Natural Language Processing (NLP) [65]. In this work, we do not explore the possibility of utilizing our benchmark for model training. Our primary goal in this paper is to provide a large-scale evaluation benchmark that addresses the issue of data leakage in current multimodal evaluation benchmarks and offers evaluation results across multiple subtasks, scenarios, question types and styles. Nevertheless, considering that Dysca has the capability to synthesize high-resolution and unlimited amounts of annotated multimodal data, we believe that Dysca also holds potential as a training data synthesis tool for LVLMs.
Reproducibility and Licence. All the experiments are built on 8 * RTX 4090. All the data and the code for generation and evaluation are released in https://github.com/Benchmark-Dysca/Dysca. The licence of Dysca is "CreativeML Open RAIL++-M", which follows the licence set by the Stable Diffusion XL.
Ethical Concerns. Our Dysca leverages the Stable Diffusion XL [36] to generate images. In order to prevent the model generating unsafe images, e.g., NSFW and offensive images, lots of efforts have been made. First, we use the safety checker [66] to post filter the unsafe images. With the unsafe image that is recognized by the safety checker, the model’s output will be a blank image. Besides, we manually exclude the specific styles or the word that may trigger the unsafe images generation from the Metadata $M$. We believe that our Dysca involves fewer ethical concerns.
@@ B.2 The Stability of Dysca
Figure 6: The tendency of model’s overall performance under clean scenario with different scale of evaluation data.
In this section, we focus on examining the stability of Dysca. We partition Dysca into 11 different scales: 1%, 10%, 20%, 30%, 40%, 50%, 60%, 70%, 80%, 90% and 100%. We compute the evaluation scores using each of these data scales. The score is calculated as the sum of scores obtained from multiple-choice, true-or-false and free-form questions. As can be seen in Fig. 6, when the evaluation data scale is less than 30% of Dysca (i.e., less than 46.8K samples), the evaluation score show significant fluctuations. When the data scale exceeds 40%, we obtain the stable results, reflecting current scale of Dysca achieves the stable and reliable evaluation results. Although 40% evaluation scale of Dysca has achieved stable scores, Dysca aims to provide more than just stable rankings, but also draws on massive amounts of data to provide in-depth feedback across different image styles and perceptual subtasks.
Appendix C The Metadata ($M$) of Dysca
@@
In order to ensure the quality and stability of the generated images, we carefully select the source material. First, for each perceptual subtask, we collect rich annotation material as described in Section 3.2. However, the metadata composed of these raw annotations is not always usable. On the one hand, some of the content is polysemous, which can easily be misinterpreted by the model’s when generating images. On the other hand, there are backgrounds or artistic styles (e.g., “Pokemon Style", “architectural style", etc.) that negatively affect the quality of the image and do not accurately generate the desired content. In order to test the usability of these source materials, we went through several small-scale pre-generations covering all the source materials. After careful selection, we retain the clips that consistently produced high-quality images. The detailed information of the source materials are shown in Tab. 11.
@@
Object We took the annotations from the MSCOCO [50] dataset after removing people, animals, and plants, and added some additional common objects. 80
@@
Appendix D Scenarios Details
D.1 Print Attack Scenario
Followed by the settings in [35], we add the attack text on the images. Consider that the image resolution in our Dysca is much higher than the one in [35], we increase more font form in terms of font position and font orientation. Fig. 7 to Fig. 11 shows the detailed examples.
@@ D.2 Corrupted Scenario
Appendix E More Examples of Dysca
@@
Appendix F Data Sheet
We follow the documentation frameworks provided by Gebru et al. [67].
F.1 Motivation
@@
  • It is funded by the Key Laboratory of AI Safety of CAS, Institute of Computing Technology.
F.2 Composition
@@ F.3 Collection Process
F.4 Preprocessing/cleaning/labeling
@@
  • Yes. We leverage the off-the-shelf models, i.e., PP-OCRv3 [56] and CLIP-L-14 [16], to clean the data. PP-OCRv3 [56] is leveraged as the filter to exclude the failure image that TextDiffusion2 [58] generates the wrong text on the image. For the other images, we use CLIP-L-14 [16] to filter out the images with low text-image consistency.

### v3

@@ B.1 General Discussion
Synthesis Data for Training / Fine-tuning. The use of synthetic data for model training / fine-tuning has been adopted in the field of Natural Language Processing (NLP) (Meta, 2024). In this work, we do not explore the possibility of utilizing our benchmark for model training. Our primary goal in this paper is to provide a large-scale evaluation benchmark that addresses the issue of data leakage in current multimodal evaluation benchmarks and offers evaluation results across multiple subtasks, scenarios, question types and styles. Nevertheless, considering that Dysca has the capability to synthesize high-resolution and unlimited amounts of annotated multimodal data, we believe that Dysca also holds potential as a training data synthesis tool for LVLMs.
Reproducibility and License. All the experiments are built on 8 * RTX 4090. All the data and the code for generation and evaluation are released at https://github.com/Robin-WZQ/Dysca. The license of Dysca is “CreativeML Open RAIL++-M”, which follows the license set by the Stable Diffusion XL.
Ethical Concerns. Our Dysca leverages the Stable Diffusion XL (Podell et al., 2023) to generate images. In order to prevent the model generating unsafe images, e.g., NSFW and offensive images, lots of efforts have been made. First, we use the safety checker (Rando et al., 2022) to post filter the unsafe images. The safety checker is a post-processor deployed by model developers (Podell et al., 2023) to prevent the generation of NSFW images. With the unsafe image that is recognized by the safety checker, the model’s output will be a blank image. Besides, we manually exclude the specific styles or the word that may trigger the unsafe images generation from the Metadata $M$. To validate the safety of Dysca, we utilize the NudeNet detector (Bedapudi, 2022) with with a threshold of 0.8 to identify images that contain NSFW content. NudeNet is the widely used automatic tool for detecting NSFW images which is adopted by (Gandikota et al., 2023; Lu et al., 2024b). With 77847 images (clean scenario) in our Dysca, only 5 images are classified as NSFW (0.006%). We check the 5 images and find they are all false positive samples. This indicates that our Dysca contains a minimal amount of such unsafety images.
@@ B.2 The Stability of Dysca
Figure 9: The tendency of 10 model’s overall performance under clean scenario with different scale of evaluation data.
In this section, we focus on examining the stability of Dysca. We partition Dysca into 11 different scales: 1%, 10%, 20%, 30%, 40%, 50%, 60%, 70%, 80%, 90% and 100%. We randomly select 10 models and compute their evaluation scores using each of these data scales. The score is calculated as the sum of scores obtained from multiple-choice, true-or-false and free-form questions. As can be seen in Fig. 9, when the evaluation data scale is less than 30% of Dysca (i.e., less than 46.8K samples), the evaluation score show significant fluctuations. When the data scale exceeds 40%, we obtain the stable results, reflecting current scale of Dysca achieves the stable and reliable evaluation results. Although 40% evaluation scale of Dysca has achieved stable scores, Dysca aims to provide more than just stable rankings, but also draws on massive amounts of data to provide in-depth feedback across different image styles and perceptual subtasks.
Appendix C The Correctness of Dysca
Since our generative pipeline relies on text-to-image models, the quality of the Dysca dataset is inherently influenced by the generation performance of the model. To address this issue, we have employed extensive data cleaning techniques to ensure the correctness of the dataset. Furthermore, to validate this, we randomly sample 7.7K images (10% of our data) and manually select images that contain wrong content or potentially lead to incorrect responses. In the end, a total of 167 images (2.2%) are filtered. Compared to previous but wide spread benchmarks, the incorrect data ratio of Dysca is less (Northcutt et al., 2021). For example, the ImageNet dataset (Deng et al., 2009) has at least 2,916 errors in its validation set, with an error rate of 6%; the QuickDraw dataset (Ha & Eck, 2018) has at least 5 million errors, with an error rate of approximately 10%. Subsequently, we compute the scores with and without these incorrect images, and the results are presented below.
Table 9: The scores with and without these incorrect images.
Model Before After
minigpt4-vicuna-7b 41.38 41.24
minigpt4-vicuna-13b 50.17 50.78
minigpt4-llama2 56.61 56.82
minigpt-v2 58.46 58.43
blip2-flan-t5-xl 65.30 65.32
blip2-opt-3b 39.54 39.80
blip2-opt-7b 39.55 39.58
instructblip-vicuna-7b 67.54 67.89
instructblip-vicuna-13b 64.89 65.02
instructblip-flan-t5-xl 66.54 66.73
instructblip-flan-t5-xxl 68.65 68.74
llava-1.5-7b 51.27 51.44
llava-1.5-13b 59.23 59.46
otter 54.90 54.75
shikra-7b 62.24 62.35
internlm-xcomposer-vl-7b 71.40 71.55
internlm-xcomposer2-vl-7b 79.13 79.23
qwen-vl-chat 62.18 62.26
emu2-chat 63.64 63.67
glm-4v-9b 82.09 82.13
minicpm-llama2-v2.5 78.75 79.02
yi-vl 75.71 76.32
mplug-owl2 74.09 74.43
phi-3-vision 73.23 73.57
[…]
Gemini-1.5-Pro-Vision-latest 77.79 77.99
As observed, the score difference is negligible, making the automatic evaluation result reliable. We believe that with the development of deep generative models, Dysca will serve as an evolving benchmark with improving quality.
Appendix D Blind Experiment
To demonstrate that Dysca has minimal data leakage issues, we conduct the "Blind" experiment where only textual questions are provided to LVLMs. We present the evaluation results of LLaVa-1.5-7B  (Liu et al., 2023b) on Dysca and the six benchmarks, including MMMU  (Yue et al., 2024), MMB  (Liu et al., 2023c), ScienceQA  (Lu et al., 2022), AI2D  (Kembhavi et al., 2016), Seed  (Li et al., 2023c) and MathVista  (Lu et al., 2024a).
Table 10: Scores of LLaVa under the blind setting across 7 benchmarks.
Model MMMU MMB ScienceQA AI2D SEED MathVista Dysca
Random Choice 22.1 0 24.2 23.8 24.3 17.9 37.5
LLaVA-1.5-7B 29.9 19.5 64.1 48.7 37.5 20.3 38.7
As shown in Tab. 10, LLAVA significantly outperforms random selection when only text questions are provided in the other 6 benchmarks. In contrast, Dysca achieves results closest to random selection, indicating that our work has a limited data leakage issue.
Appendix E The Metadata ($M$) of Dysca
@@
In order to ensure the quality and stability of the generated images, we carefully select the source material. First, for each perceptual subtask, we collect rich annotation material as described in Section 3.2. However, the metadata composed of these raw annotations is not always usable. On the one hand, some of the content is polysemous, which can easily be misinterpreted by the model’s when generating images. On the other hand, there are backgrounds or artistic styles (e.g., “Pokemon Style”, “architectural style”, etc.) that negatively affect the quality of the image and do not accurately generate the desired content. In order to test the usability of these source materials, we went through several small-scale pre-generations covering all the source materials. After careful selection, we retain the clips that consistently produced high-quality images. The detailed information of the source materials are shown in Tab. 11.
@@
Object We took the annotations from the MSCOCO (Lin et al., 2014) dataset after removing people, animals, and plants, and added some additional common objects. 80
@@
Appendix F Scenarios Details
F.1 Print Attack Scenario
Followed by the settings in (Cheng et al., 2024), we add the attack text on the images. Consider that the image resolution in our Dysca is much higher than the one in (Cheng et al., 2024), we increase more font form in terms of font position and font orientation. Fig. 10 to Fig. 14 shows the detailed examples.
@@ F.2 Corruption Scenario
Appendix G Hard Samples
In this section, we provide hard samples that LVLMs are likely to answer incorrectly, as shown in Fig. 16 to Fig. 18.
Figure 16: Hard samples.
Figure 17: Hard samples.
Figure 18: Hard samples.
Appendix H More Examples of Dysca
@@
Appendix I JSON structure for evaluation
Here, We present two example JSON structures generated by Dysca for evaluation in Listing 1.
Listing 1: Example JSON structure for evaluation
"id": 1,
"images": [
"images/1.png"
"prompt": "face portrait of a 25 years old female, Indian race, calm and placid, on Mars, retro arcade style, 8-bit, pixelated, vibrant, classic video game, old school gaming, reminiscent of 80s and 90s arcade games",
"negative_prompt": "modern, ultra-high resolution, photorealistic, 3D",
[…]
"question": "What is the background in the picture?",
"answer": "on Mars",
"question_type": "multi choice",
"task": "recognition",
"options": [
"on the lawn",
"under the water background",
"sunset background",
"on Mars"
"question_majority": "background",
"granularity": "coarse"
"id": 2,
"images": [
"images/2.png"
"prompt": "1 white color cat, on Mars, thick layered papercut art of, deep 3D, volumetric, dimensional, depth, thick paper, high stack, heavy texture, tangible layers",
"negative_prompt": "2D, flat, thin paper, low stack, smooth texture, painting, drawing, photo, deformed",
[…]
"question": "Is there a cat in the picture?",
"answer": "True",
"question_type": "true or false",
"task": "recognition",
"options": [
"False",
"True"
"question_majority": "animal",
"granularity": "fine"
Appendix J Data Sheet
We follow the documentation frameworks provided by Gebru et al. Gebru et al. (2021).
J.1 Motivation
@@ Appendix I JSON structure for evaluation
  • This work is partially supported by Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680202), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No. SYG202325), and Youth Innovation Promotion Association CAS.
J.2 Composition
@@
J.4 Preprocessing/cleaning/labeling
@@
  • Yes. We leverage the off-the-shelf models, i.e., PP-OCRv3 Li et al. (2022) and CLIP-L-14 Radford et al. (2021), to clean the data. PP-OCRv3 Li et al. (2022) is leveraged as the filter to exclude the failure image that TextDiffusion2 Chen et al. (2023a) generates the wrong text on the image. For the other images, we use CLIP-L-14 Radford et al. (2021) to filter out the images with low text-image consistency.